Use α = 0.01 and conduct a goodness of fit test using 5 classes to see whether the following sample appears to have been selected from a normal probability distribution.

(Note: x = 70 and s = 18.)

55 86 94 58 55 95 55 52 69 95 90 65 87 50 56
55 57 98 58 79 92 62 59 88 40

Find the value of the test statistic.

Find the p-value. (Round your answer to four decimal places.)

p-value =
Answer #1

Goodness of Fit test(Chi square):

Ho: To test the data is distributed normally and sampled from normal distribution

Ha: To test the data is not distributed normally and sample is not normal

Significant level: 0.01

The degree of freedom: n-1

n=25

DF=25-1=24

Significant value: chi ^2_{df,alpha }= chi ^2_{24,0.01}=42.98

Distribution Plot Chi-Square, df 24 0.06 0.05 0.04- 0.03 0.02 0.01 0.01 0.00 42.98

The test statistic:

chi ^2= rac{sum (O_i-E_i)^2}{E_i}

O- Observation, E- Expected value(Mean)

X Mean (X-mean)^2/Mean
55 70 3.214285714
86 70 3.657142857
94 70 8.228571429
58 70 2.057142857
55 70 3.214285714
95 70 8.928571429
55 70 3.214285714
52 70 4.628571429
69 70 0.014285714
95 70 8.928571429
90 70 5.714285714
65 70 0.357142857
87 70 4.128571429
50 70 5.714285714
56 70 2.8
55 70 3.214285714
57 70 2.414285714
98 70 11.2
58 70 2.057142857
79 70 1.157142857
92 70 6.914285714
62 70 0.914285714
59 70 1.728571429
88 70 4.628571429
40 70 12.85714286
Sum 111.8897

Σ(0-E,)2 Ei 2 111.8857

P-value: 0.0000

The test statistic is significant and rejects H0, that the sample is not normally distributed.